Circular No. 7978

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2002 Q2 (LINEAR) AND C/2002 Q3 (LINEAR) Further to IAUC 7966, Z. Sekanina writes that he has calculated an

improved set of fragmentation parameters based on 73 differential astrometric observations taken between Aug. 26 and Sept. 13 (MPC 46490–46491). The derived separation time of C/2002 Q3 from C/2002 Q2 is 1970.2 (± 1.2 yr), when the parent comet was at a distance of 56 (± 1) AU from the sun, the separation velocity was 0.18 (± 0.01) m/s, and C/2002 Q3 (component A) was subjected to a relative deceleration of 34.8 (± 1.3) units of 10^{−5} solar gravitational acceleration. The predicted separation ephemeris differs very little from that on IAUC 7966. A fragmentation sequence for the components of C/2002 Q3 itself has been determined, based on the astrometry from the period 2002 Sept. 4–14. It is found that (i) nucleus A, the original, eastern component, is the primary, (ii) the component labelled C on Sept. 10 and 14 is identical with C on Sept. 4–5, and (iii) the new, southern condensation from Sept. 10 and 14, called D, is different from B (the observations attributed to B on MPEC 2002-R54 apparently not referring to what was observed on Sept. 4–5). [The offsets of B from A on Sept. 4–5 show irreconcilable discrepancies (in excess of 3′′) when attempts are made to link them with the offsets of the C component on Sept. 10 and 14.] It appears that C separated from A before 1980, a satisfactory solution indicating a fragmentation time in 1973 (± 4 yr), at ∼ 52 AU from the sun, a separation velocity of ∼ 0.1 m/s, and a relative deceleration of ∼ 6 units of 10^{−5} solar gravitational acceleration. One cannot exclude the possibility that C separated directly from C/2002 Q2 about the time, or even before, component A did. The only satisfactory solution for D suggests that it broke off from A on 2002 July 5 ± 2 days (at ∼ 1.5 AU from the sun) with a relative velocity of 1.3 m/s and has been subjected to a deceleration of 12 ± 1 units. The predicted offsets of D from A on Sept. 4–5 are −2′′.4 in α and −8′′.9 in δ. The history of nucleus B remains unknown. Predicted separation distances (and position angles) for components C and D relative to A (0^{h} TT): 2002 Sept. 18, 3′′.1 (277^{o}) and 5′′.0 (237^{o}); 23, 1′′.5 (245^{o}) and 4′′.7 (292^{o}); 28, 1′′.8 (169^{o}) and 7′′.4 (325^{o}); Oct. 3, 3′′.3 (147^{o}) and 10′′.4 (337^{o}); 8, 4′′.7 (139^{o}) and 12′′.6 (343^{o}); 13, 5′′.7 (134^{o}) and 14′′.1 (346^{o}); 18, 6′′.4 (130^{o}) and 15′′.1 (348^{o}); 23, 7′′.0 (127^{o}) and 15′′.7 (350^{o}).

V4742 SAGITTARII CCD V magnitudes by J. D. West, Mulvane, KS: Sept. 22.074 UT,

8.77; 23.064, 9.54; 24.083, 9.22. Also: Sept. 22.084, B = 9.85.

2002 September 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green